a uniform rod of length 5m and weight 20N is supported at the 1m mark and at the 4m mark. A painter of weight 750N stands at the 3m mark. Calculate the reaction at each support.
1
Expert's answer
2012-10-10T09:36:15-0400
2 * 16 + 2 * 750 - 3 * F2 - 0.5 * 4 = 0
1 * 750 + 2 * 16 - 3 * F1 - 0.5 * 4 = 0

F1 = (750 + 32 - 2) / 3 = 26 N
F2 = (1500 + 32 - 2) / 3 = 510 N